I don't understand how to solve and check the quadratic equation below by taking the square root of both sides.
(a+5)^2=9
thank you

Answer by Maths68(1474)

(a+5)^2=9
(a+5)^2=3^2
Take squre root of both sides
sqrt((a+5)^2)=sqrt(3^2)
sqrt((a+5)*(a+5))=sqrt(3*3)
a+5=3
a=3-5
a=-2
Check
(a+5)^2=9
Put the value of a in above equation
(-2+5)^2=9
(3)^2=9
3*3=9
9=9